Trigonometric identities


Solve and prove trigonometric identities


1.) Cot(theta)sin(theta)cos(theta)=1-sin²(theta)


2.) Sin²(theta)sec²(theta)-1


3.) Tan(theta)+cot(theta)= csc(theta)sec(theta)


4.) 1+cot(theta)= csc(theta)(cos(theta) + sin(theta))




1
Expert's answer
2021-02-10T01:10:51-0500

"1.)\cot(\theta)\sin(\theta)\cos(\theta)=1-\sin^2(\theta)"

"\frac{cos(\theta)}{sin(\theta)}(\theta)\sin(\theta)\cos(\theta)=1-\sin^2(\theta)"

"\cos^2(\theta)= 1-\sin^2(\theta)"

"\cos^2(\theta)+\sin^2(\theta)=1"

"1=1\ \blacksquare"


"2.)\sin^2(\theta)\sec²(\theta)-1="

"\sin^2(\theta)\frac{1}{cos²(\theta)}-1="

"\frac{1-\cos^2(\theta)}{\cos^2(\theta)}-1=\frac{1}{\cos^2(\theta)}-2="

"\sec^2{(\theta)}-2"


"3.)\tan(\theta)+\cot(\theta)= \csc(\theta)\sec(\theta)"

"\frac{\sin(\theta)}{\cos(\theta)}+\frac{\cos(\theta)}{\sin(\theta)}= \csc(\theta)\sec(\theta)"

"\frac{\sin^2(\theta)+sin^2(\theta)}{\cos(\theta)\sin(\theta)}= \csc(\theta)\sec(\theta)"

"\frac{1}{\cos(\theta)\sin(\theta)}= \csc(\theta)\sec(\theta)"

"\frac{1}{\cos(\theta)}\frac{1}{\sin(\theta)}= \csc(\theta)\sec(\theta)"

"\sec(\theta)\csc(\theta)=\csc(\theta)\sec(\theta)"

"\csc(\theta)\sec(\theta)=\csc(\theta)\sec(\theta)\ \blacksquare"


"4.)1+\cot(\theta)= \csc(\theta)(cos(\theta) + sin(\theta))"

"1+\cot(\theta)= \frac{1}{sin(\theta)}(cos(\theta) + sin(\theta))"

"1+\cot(\theta)= \frac{\cos(\theta)}{\sin(\theta)}+1"

"1+\cot(\theta)= \cot(\theta)+1"

"1+\cot(\theta)=1+\cot(\theta)\ \blacksquare"